Circular No. 8064

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ab IN UGC 4930 Further to IAUC 8062, T. Boles reports the discovery, on unfiltered

CCD images taken on Feb. 1.971 and 2.856 UT, of an apparent supernova (mag 17.8) located at α = 9^{h}19^{m}13^{s}.26, δ = +48^{o}58′09′′.1, which is ≈ 9′′.3 east and 4′′.2 north of the center of UGC 4930. The new object is not present on Boles’ images from 2002 Dec. 29 (limiting mag 19.0) or Apr. 11, and it is not present on Palomar Sky Survey red or blue plates.

SUPERNOVA 2003ac IN IC 3203 Further to IAUC 8063, B. Swift and W. Li report the LOTOSS dis-

covery of an apparent supernova (mag ∼ 18.9) on unfiltered KAIT images taken on Jan. 7.5, 20.4, Feb. 2.5, and 3.4 UT. SN 2003ac is located at α = 12^{h}21^{m}47^{s}.35, δ = +25^{o}52′33′′.1 (equinox 2000.0), which is 23′′.4 east and 30′′.9 south of the nucleus of IC 3203. A KAIT image taken on 2002 May 17.3 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2003aa IN NGC 3367 Further to IAUC 8063, T. Matheson, P. Challis, and R. Kirshner re-

port that a spectrum (range 370–750 nm) of SN 2003aa (cf. IAUC 8063), obtained by P. Berlind with the 1.5-m reflector on Feb. 2.31 UT, shows it to be a type-Ic supernova. The spectrum is similar to that of SN 1994I, four days before maximum (Filippenko et al. 1995, Ap.J. 450, L11).

XTE J0119–731 R. Corbet, Goddard Space Flight Center (GSFC) and Universities

Space Research Association; C. B. Markwardt, GSFC and University of Maryland; F. E. Marshall, GSFC; M. J. Coe and W. R. T. Edge, University of Southampton; and S. Laycock, Center for Astrophysics, report that RXTE Proportional Counter Array observations (2–10 keV; total exposure time 13740 s) on Jan. 5 detected a transient x-ray pulsar with intensity ≈ 1.3 counts s^{−1} PCU^{−1}, or ∼ 0.625 mCrab, and a period of 2.1652 ± 0.0001 s, in the direction of the Small Magellanic Cloud (SMC): “If this represents the rotation period of a neutron star, then it is the fastest known transient pulsar in the SMC. The source position, determined from slews over this region, is α = 1^{h}19^{m}51^{s}, δ = −73^{o}11′.8 (equinox 2000.0; estimated errors 2′ in α and 3′.7 in δ). The pulsar was not detected in observations made in the same pointing direction on 2002 Dec. 13 and 2003 Jan. 17.”

2003 February 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green